OpenStudy (anonymous):

Micah can fill a box with books in 15 minutes. While Micah fills the box, Sydney takes the books out puts them on a shelf. With Sydney taking books out of the box, it takes 20 minutes for Micah to fill the box. Which of the following can be used to determine the amount of time it takes for Sydney to empty the box if Micah is not adding books? a. 1 over 15 minus 1 over 20 equals 1 over x b. 1 over 20 minus 1 over x equals 1 over 15 c. 1 over 15 minus 1 over x equals 1 over 20 d. 1 over 15 minus 1 over x equals x over 20

OpenStudy (mathstudent55):

Filling the box is doing 1 complete job. In 1 minute, Micah does 1/15 of the job In 1 minute, Sydney undoes 1/x of the job In 1 minute, Micah and Sydney do 1/20 of the job 1/15 - 1/x = 1/20

OpenStudy (tkhunny):

Standard Problem -- Working Together x = The Rate at which Sydney Helps 1/15 + 1/x = 1/(Something less than 15) NonStandard Problem -- Not helping x = The Rate at which Sydney Helps 1/15 - 1/x = 1/(Something greater than 15) What say you?